OpenStudy (anonymous):

@zepdrix Find the slope of the tangent line to the graph of 4x^2 y- pi cos y=5 pi at the point (1, pi).

zepdrix (zepdrix):

\[\large 4x^2y-\pi \cos y=5\pi\] Hmm ok this will work out similar to the normal line problem we did earlier. We won't have to worry about the negative reciprocal nonsense this time though.

OpenStudy (anonymous):

lol that's better

zepdrix (zepdrix):

Have you tried taking the derivative yet?

OpenStudy (anonymous):

No, not yet! I got a little confused cause there are so many different variables

zepdrix (zepdrix):

For the first term, we'll apply the product rule.\[\large \left(4x^2y\right)' \qquad=\qquad (4x^2)'y+4x^2(y)'\]

OpenStudy (anonymous):

Okay

zepdrix (zepdrix):

So the first term when we take it's derivative is giving us,\[\large 8xy+4x^2y'\] Uh oh, you're poppin out the "okay" 's. lol :) confused i'll bet

OpenStudy (anonymous):

No I'm good, I'm working it on my paper as well so I understand! So far so good!

OpenStudy (anonymous):

You're seriously so awesome at teaching me, you don't even know!

zepdrix (zepdrix):

In the next term, that pi is just a constant. So let's not worry too much about it. Derivative of -cosine is umm positive sine right?\[\large \left(-\pi \cos y\right)' \qquad=\qquad \pi (\sin y)y'\] And we need to recall that any time we take the derivative of something containing y, we plop a y' onto the end of it. I put brackets around the sin y just so there's no confusion that the y' is NOT inside of the sine function with y.

OpenStudy (anonymous):

So far so good!

zepdrix (zepdrix):

The right side is constant. Derivative of a constant? :O

OpenStudy (anonymous):

Is 0!

zepdrix (zepdrix):

Ok sounds good! :) So this is what we have so far,\[\large 8xy+4x^2y'+\pi(\sin y)y'=0\]

zepdrix (zepdrix):

I'm glad we did those other two problems BEFORE this one. Because we'll be using techniques from both of those problems. :) So it's a nice lead-in to this one.

OpenStudy (anonymous):

Lol yay! :)

zepdrix (zepdrix):

Let's get the y' terms on one side. So let's subtract 8xy from each side,\[\large 4x^2y'+\pi(\sin y)y'=-8xy\]Then we'll factor a y' out of each term,\[\large y'\left(4x^2+\pi \sin y\right)=-8xy\]Then divide by some stuff.

OpenStudy (anonymous):

So I would I just put in the (1, pi)?

zepdrix (zepdrix):

\[\large y_t=mx+b\] Our particular slope \(\large m\) is given by,\[\large y'(1,\pi)=m\] Yes good :)

OpenStudy (anonymous):

So the answer is about 2 pi! Wow so helpful! :D

zepdrix (zepdrix):

haha `about` 2pi? XD

OpenStudy (anonymous):

Yeah, I got a decimal lol, not exactly 2 pi

zepdrix (zepdrix):

Hmm weird, I got -2pi.

OpenStudy (anonymous):

I mean negative! But did you get that exactly?

zepdrix (zepdrix):

\[\large y'(x,y)=\frac{-8xy}{4x^2+\pi \sin y}\] \[\large y'(1,\pi)=\frac{-8(1)(\pi)}{4(1)^2+\cancel{\pi \sin \pi}}\]

zepdrix (zepdrix):

\[\large y'(1,\pi)=\frac{-8\pi}{4}\]Hmm ya I got exactly :o

zepdrix (zepdrix):

Sine of pi is zero, did you punch it into your calculator in degree mode or something silly like that?

OpenStudy (anonymous):

Lol cause I put it in the calculator, but the answer doesn't come out as pi you have to convert it unless there is some button I don't know about. But doing it by hand like you did helps too!

zepdrix (zepdrix):

oh i see

OpenStudy (anonymous):

Okay, I just have a few more I promise! You are really helping me do good on my test tomorrow! I thought I would have to be up for hours trying to figure these out!

zepdrix (zepdrix):

Ok post another one, then imma take a lil math break after that :D Lots of people on here to help you though, don't worry!

OpenStudy (anonymous):

No one seems to respond or know how to do them lol